A. If a firm produce and sells q units of a particular product in the market. The total cost of producing q units is given by the function C(q) = 5q + 350 while the price firm received in the market is given by the function D(q) = 50 - q . Find



a. The profit function



b. The break-even value of q



c. The value of q that result in loss

1
Expert's answer
2022-03-23T18:32:35-0400

a) Profit (P) is defined as total revenue (TR) minus Total Cost (TC). Therefore,

"P = TR-TC"

"P =50-q-(5q+350)"

"=50+350-q-5q"

Thus;

"P = 400-6q"

b) The breakeven point is when revenue equals cost or when the profit is zero. Therefore, break even point at value q;

"0=400-6q"

"-400=-6q"

"q=66.67"

The break even value is 66.67 units of q.

c) loss is made when the revenue from sales is not enough to cover all the costs of production.

Therefore;

"50-q=-5q+350"

"50-350=-5q+q"

"-300=-4q"

q=75